# Finite-time blowup in a five- and higher-dimensional chemotaxis model with indirect signal production

### v2

@@
Abstract.
This paper is concerned with a three-component chemotaxis model accounting for indirect signal production, reading as $u_{t}=\nabla\cdot(\nabla u-u\nabla v)$, $v_{t}=\Delta v-v+w$ and $0=\Delta w-w+u$, posed on a ball in $\mathbb{R}^{n}$ with $n\geq 5$, subject to homogeneous Neumann boundary conditions.
It is suggested that the system has a four-dimensional critical mass phenomenon concerning blowup in finite or infinite time according to the work of Fujie and Senba [J. Differential Equations, 263 (2017), 88–148; 266 (2019), 942–976]. In this paper, we prove that for any prescribed mass $m>0$, there exist radially symmetric and positive initial data $(u_{0},v_{0})\in C^{0}(\overline{\Omega})\times C^{2}(\overline{\Omega})$ with $\int_{\Omega}u_{0}=m$ such that the corresponding solutions blow up in finite time.
Key words and phrases:
chemotaxis; indirect signal production; finite-time blowup
2020 Mathematics Subject Classification:
35B44; 35K51; 35Q92; 92C17
Supported in part by National Natural Science Foundation of China (No. 12271092, No. 11671079) and the Jiangsu Provincial Scientific Research Center of Applied Mathematics (No. BK20233002).
1. Introduction
@@
$\begin{cases}u_{t}=\Delta u-\nabla\cdot(u\nabla v),&x\in\Omega,t>0,\\ v_{t}=\Delta v-v+w,&x\in\Omega,t>0,\\ 0=\Delta w-w+u,&x\in\Omega,t>0,\\ \partial_{\nu}u=\partial_{\nu}v=\partial_{\nu}w=0,&x\in\partial\Omega,t>0,\\ (u(\cdot,0),v(\cdot,0))=(u_{0},v_{0}),&x\in\Omega,\end{cases}$ (1.1)
@@
Among the chemotaxis models that describe the oriented movement of organisms in response to chemical stimuli, the minimal chemotaxis system [35]
$\begin{cases}u_{t}=\nabla\cdot(\nabla u-u\nabla v),&x\in\Omega,t>0,\\ v_{t}=\Delta v-v+u,&x\in\Omega,t>0,\end{cases}$ (1.2)
has been extensively studied. This system was originally proposed by Keller and Segel  [19] to investigate the aggregation of cells. The density $u$ of cells is influenced by their motion, which consists of random walk and chemotactic movement along the gradient of chemical concentration $v$. As is well-known, the system exhibits a competitive mechanism between diffusion and aggregation induced by chemotaxis where blowup—the extreme form of aggregation—can occur in finite time under certain conditions:
  • In one dimension, diffusion dominates aggregation, and all solutions of (1.2) are globally bounded [37].
@@
We highlight some related work on (1.2) for interested readers, regarding various aspects: chemotactic collapse [33], asymptotic behaviors under assumptions of uniform-in-time boundedness [9], subcritical mass [44] in two dimensions or smallness conditions [3] in higher dimensions, energy criterion for finite-time blowup [29], blowup rate [30] and blowup profile [48].
Recently, chemotaxis models with indirect signal production, often expressed in the following typical form
$\begin{cases}u_{t}=\Delta u-\nabla\cdot(u\nabla v)+f(u,w),&x\in\Omega,t>0,\\ v_{t}=\Delta v-v+w,&x\in\Omega,t>0,\\ w_{t}=\tau\Delta w-w+u,&x\in\Omega,t>0,\end{cases}$ (1.3)
@@
When the signal producer is static, i.e., $\tau=0$, the indirect attractive mechanism has an explosion-suppressing (or boundedness-supporting) effect. Hu and Tao [16] showed that quadratic logistic dampening $f(u,w)=\mu u(1-u)$ to any weak extent $\mu>0$ can prevent blowup in a three-dimensional indirect chemotaxis model with growth (1.3), In contrast, Winkler [46] established boundedness in the three- and higher-dimensional parabolic-parabolic chemotaxis system with a strong logistic source in the sense of $\mu>\mu_{\star}$ for some $\mu_{\star}>0$. For the mass-preserving system (1.3) with $f\equiv 0$, a novel two-dimensional critical mass phenomenon has been identified, concerning infinite-time blowup as documented by Tao and Winkler [42] and Laurençot [21]. For chemotaxis models (1.3) with phenotype switching $f=-u+w$, the occurrence of two-dimensional infinite-time blowup depends on the initial total mass $\int_{\Omega}u_{0}+\int_{\Omega}v_{0}$ [23].
@@
For fully parabolic indirect-chemotaxis models ($\tau=1$) without a source term $f(u,w)\equiv 0$, the indirect attractant mechanism can also reduce explosion-supporting potential. Fujie and Senba [10, 11] established global boundedness in physical spaces $n\leq 3$ and identified a four-dimensional critical mass phenomenon for classical solutions of (1.3). Specifically, they found that
@@
  • Under symmetry assumptions, the solutions exists globally and remain bounded for nonnegative initial data $(u_{0},v_{0},w_{0})$ with $\int_{\Omega}u_{0}<64\pi^{2}$. Without radial symmetry, the same conclusion holds for (1.3) under mixed boundary conditions, where $\partial_{\nu}u-u\partial_{\nu}v=v=w=0$ on $\partial\Omega$ for $t>0$.
@@
  • There exist initial data satisfying $\int_{\Omega}u_{0}\in(64\pi^{2},\infty)\setminus 64\pi^{2}\mathbb{N}$ such that the solutions of (1.3) under mixed boundary conditions blow up in finite or infinite time.
@@
Beyond four dimensions, Mao and Li showed that there exist initial data such that the solutions of the system (1.3) blow up in finite or infinite time [28] with symmetry assumptions. We emphasize both critical dimension $4$ and critical mass $64\pi^{2}$ of (1.3) regarding blowup are significantly larger compared to those of the classical chemotaxis model (1.2). This increase is due to the strong parabolic smoothing effect. Chen and Xiang [5] observed a similar positive effect on boundedness in certain indirect chemotaxis-haptotaxis model. Additionally, taxis-driven aggregation, particularly the spontaneous emergence of large densities during the intermediate transition triggered by a large phenotype switching rate, has been explored in [39, 24].
@@
Concerning finite-time singularity formation, Tao and Winkler [43] considered a Jäger-Luckhaus variant [17] of the system (1.3), where the second equation was modified to $\Delta v-\int_{\Omega}w/|\Omega|+w=0$, and the source term was taken as $f=-au+bw$ for some constants $a\geq 0$ and $b\geq 0$. They showed that there exist radially symmetric initial data such that solutions of the variant, posed on a ball in five and higher dimensions, blow up in finite time. This was deduced by robust comparison methods, which are applicable to cooperative systems involving volume-filling [38] and phenotype switching [27]. In this regard, we focus on detecting finite-time blowup for (1.1), a Nagai-type variant [32] of (1.3). Here, we adopt the assumption that the change of $w$ is stationary over time [18]. However, the comparison arguments used in previous works may not be directly applicable to (1.1) due to the loss of certain partial order relations, inter alia.
@@
The system (1.1) has an energy dissipation structure identified by Fujie and Senba [10]. The Lyapunov functional,
@@
along the reasonably regular solution curves of the system (1.1). Here, the dissipation rate reads as
@@
$\begin{cases}-\Delta w+w=u,&x\in\Omega,\\ \partial_{\nu}v=0,&x\in\partial\Omega.\end{cases}$ (1.4)
Invoking the above mathematical structure, we are inspired by previous results (see Remark 1.3) to derive a lower bound for $\mathcal{F}$ in terms of a sublinear power of $\mathcal{D}$ in the form of
$-\mathcal{F}\leq C(\mathcal{D}^{\theta}+1)\quad\text{with some }\theta\in(0,1)\text{ and }C>0,$ (1.5)
for a large class of radially symmetric function pairs $(u,v)=(u(r),v(r))$, which, in particular, includes the radial trajectories of the system (1.1). This result leads to a superlinear differential inequality for $\mathcal{F}$, which, in turn, implies occurrence of finite-time blowup for initial data with low energy.
@@
$0\leq u_{0}\in C^{0}(\overline{\Omega}),\quad 0\leq v_{0}\in C^{2}(\overline{\Omega})\quad\text{with }\partial_{\nu}v_{0}\equiv 0\text{ on }\partial\Omega$ (1.6)
@@
$(u_{0},v_{0})\text{ is a pair of positive and radially symmetric functions on }\overline{\Omega},$ (1.7)
@@
[…]
$\text{and }\mathcal{F}(u_{0},v_{0})<-K(m,A)\},$
@@
Concerning the existence of initial data with low energy, for any given positive initial datum with suitable regularity, we construct a family of initial data with large negative energy, which can approximate the given initial datum in an appropriate topology.
@@
Assume $\Omega=B_{R}\subset\mathbb{R}^{n}$ for some $n\geq 5$ and $R>0$. If $u_{0}\in C^{0}(\overline{\Omega})$ and $v_{0}\in C^{2}(\overline{\Omega})$ comply with (1.6) and (1.7), then one can find a family
$\{(u_{\eta},v_{\eta})\in L^{\infty}(\Omega)\times W^{2,\infty}(\Omega)\}_{\eta\in(0,\eta_{\star})}$
of function couples for some $\eta_{\star}\in(0,1)$ with properties: $(u_{\eta},v_{\eta})$ satisfying (1.6), (1.7) and $\int_{\Omega}u_{\eta}=\int_{\Omega}u_{0}$ for all $\eta\in(0,\eta_{\star})$ is such that
$\|u_{\eta}-u_{0}\|_{L^{p}(\Omega)}+\|v_{\eta}-v_{0}\|_{W^{2,2}(\Omega)}\to 0\quad\text{as }\eta\searrow 0,$
holds for all
$p\in\bigg{[}1,\frac{2n}{n+4}\bigg{)},$
but
$\mathcal{F}(u_{\eta},v_{\eta})\to-\infty\quad\text{as }\eta\searrow 0.$
In particular, for any $\varepsilon>0$, there exist $\eta_{\star\star}=\eta_{\star\star}(\varepsilon)\in(0,\eta_{\star})$, $K(m,A+\varepsilon)$ and $T(m,A+\varepsilon)$ with the property in Theorem 1.1 such that for all $\eta\in(0,\eta_{\star\star})$, $(u_{\eta},v_{\eta})\in\mathcal{B}(m,A+\varepsilon)$ defined as in (1.8), where $\|u_{0}\|_{L^{1}(\Omega)}=m$ and $A=\|v_{0}\|_{W^{2,2}(\Omega)}$.
Remark 1.3.
Our methods are built on the original ideas of Winkler [47], and are also inspired by previous works where functional inequalities of the type in (1.5) have been successfully applied to chemotaxis models in various contexts: volume-filling by Cieślak and Stinner [6, 7, 8], by Laurençot and Mizoguchi [22], by Hashira, Ishida and Yokota [13] and by Cao and Fuest [4]; the Cauchy problem by Winkler [49]; competing chemotaxis by Lankeit [20]; two species models by Li and Li [25] and by Ha and Kim[12]. As far as we are aware, this is the first work to adapt the functional inequality (1.5) to the context of indirect chemotaxis.
The primary challenge lies in deriving the functional inequality (1.5) due to explosion-suppressing effect of indirect attractant mechanisms. Mathematically, we shall obtain interior estimates $\|\Delta v\|_{L^{2}(B_{\rho})}$ in terms of the dissipation rate $\mathcal{D}$ for $\rho\in(0,R)$. To achieve this, we first establish weighted uniform estimates $\||x|^{\beta-1}v\|_{W^{1,\infty}(\Omega)}$ with some $\beta>n-2$ in Section 3. In Section 4, we introduce new techniques to handle the more complex second-order estimate.
This paper is organized as follows. In Section 2, we present some preliminaries. Section 3 is dedicated to deriving uniform-in-time pointwise estimates for the chemical concentration. In Section 4, we establish the functional inequality (1.5). Section 5 is devoted to the proof of Theorem 1.1. In Section 6, we construct a family of positive initial data with arbitrarily large negative energy thus completing the proof of Theorem 1.2.
2. Preliminaries. Energy functional and basic estimates
We begin by revisiting the results on the local existence and uniqueness of classical solutions to the system (1.1).
@@
Let $n\in\mathbb{N}$ and $\Omega\subset\mathbb{R}^{n}$ be a bounded domain with smooth boundary. Assume that $(u_{0},v_{0})$ is as in (1.6). Then there exist $T_{\max}\in(0,\infty]$ and a unique triplet $(u,v,w)$ of nonnegative functions from $C^{0}\left(\bar{\Omega}\times\left[0,T_{\max}\right)\right)\cap C^{2,1}\left(\bar{\Omega}\times\left(0,T_{\max}\right)\right)$ solving (1.1) classically in $\Omega\times\left(0,T_{\max}\right)$. Moreover $u>0$ in $\Omega\times\left(0,T_{\max}\right)$ and
$\text{if }T_{\max}<\infty,\text{ then }\|u(\cdot,t)\|_{L^{\infty}(\Omega)}\to\infty\text{ as }t\nearrow T_{\max}.$ (2.1)
If in addition $\Omega=B_{R}$ for some $R>0$, and $(u_{0},v_{0})$ is a triplet of radially symmetric functions, then $u$, $v$ and $w$ are all radially symmetric.
@@
The local existence, uniqueness and extensibility criterion of classical solutions to the system (1.1) can be established by adapting well-known fixed-point arguments in the contexts of chemotaxis models, as outlined in [15, Theorem 3.1], [41, Lemma 3.1] and [10, Section 4]. The nonnegativity follows from the maximum principle, while conservation of radial symmetry is a consequence of the uniqueness of solutions and the adequate form of the equations in (1.3). ∎
A Lyapunov functional associated with the system (1.1) was formulated in [10, Proposition 6.1].
@@
$\frac{d}{dt}\mathcal{F}(u(t),v(t))+\mathcal{D}(u(t),v(t))=0\quad\text{ for all}t\in(0,T_{\max})$ (2.2)
@@
$\mathcal{F}(u,v)=\int_{\Omega}(u\log u-uv)+\frac{1}{2}\int_{\Omega}|(-\Delta+1)v|^{2}\quad\text{and}$ (2.3)
$\mathcal{D}(u,v)=\int_{\Omega}\left(\left|\nabla v_{t}\right|^{2}+\left|v_{t}\right|^{2}\right)+\int_{\Omega}u|\nabla(\log u-v)|^{2}.$
@@
Let $n\geq 4$ and $(u,v,w)$ be a classical solution given as in Proposition 2.1. Then
$\int_{\Omega}u(x,t)=\int_{\Omega}u_{0}=\int_{\Omega}w(x,t)=:m\quad\text{for all }t\in(0,T_{\max})$ (2.4)
@@
$\int_{\Omega}v(x,t)\leq\max\left\{\int_{\Omega}v_{0},\int_{\Omega}u_{0}\right\}\quad\text{for all }t\in(0,T_{\max}).$ (2.5)
For each
$q\in\left[1,\frac{n}{n-1}\right)\quad\text{and}\quad p\in\left[1,\frac{n}{n-2}\right),$
there exists $C_{q}>0$ and $C_{p}>0$ such that
$\|w\|_{W^{1,q}(\Omega)}\leq C_{q}m\quad\text{for all }t\in(0,T_{\max})$ (2.6)
@@
$\|v_{t}\|_{L^{p}(\Omega)}+\|w\|_{L^{p}(\Omega)}+\|v\|_{W^{2,p}(\Omega)}\leq C_{p}(m+\|v_{0}\|_{W^{2,2}(\Omega)})\quad\text{for all }t\in(0,T_{\max}),$ (2.7)
respectively.

### v3

@@
Abstract
This paper is concerned with a three-component chemotaxis model accounting for indirect signal production, reading as $u_{t}=\nabla\cdot(\nabla u-u\nabla v)$, $v_{t}=\Delta v-v+w$ and $0=\Delta w-w+u$, posed in a ball of $\mathbb{R}^{n}$ with $n\geq 5$, subject to homogeneous Neumann boundary conditions. The system is a Nagai-type variant of its fully parabolic version that has a four-dimensional critical mass phenomenon concerning blowup in finite or infinite time according to the seminal works of Fujie and Senba [J. Differential Equations, 263 (2017), 88–148; 266 (2019), 942–976]. We prove that for any prescribed mass $m>0$, there exist radially symmetric and positive initial data $(u_{0},v_{0})\in C^{0}(\overline{\Omega})\times C^{2}(\overline{\Omega})$ with $\int_{\Omega}u_{0}=m$ such that the corresponding solutions blow up in finite time.
keywords:
chemotaxis, indirect signal production, finite-time blowup
pacs:
MSC Classification]Primary 35B44; Secondary 35K51, 35Q92, 92C17
1 Introduction
@@
$\begin{cases}u_{t}=\Delta u-\nabla\cdot(u\nabla v),&x\in\Omega,t>0,\\ v_{t}=\Delta v-v+w,&x\in\Omega,t>0,\\ 0=\Delta w-w+u,&x\in\Omega,t>0,\\ \partial_{\nu}u=\partial_{\nu}v=\partial_{\nu}w=0,&x\in\partial\Omega,t>0,\\ (u(\cdot,0),v(\cdot,0))=(u_{0},v_{0}),&x\in\Omega,\end{cases}$ (1)
@@
Among the chemotaxis models that describe the oriented movement of organisms in response to chemical stimuli, the minimal chemotaxis system
$\begin{cases}u_{t}=\nabla\cdot(\nabla u-u\nabla v),&x\in\Omega,t>0,\\ v_{t}=\Delta v-v+u,&x\in\Omega,t>0,\end{cases}$ (2)
has been extensively studied. This system was originally proposed by Keller and Segel [1] to investigate the aggregation of cells. The density $u$ of cells is influenced by their motion, which consists of random walk and chemotactic movement along the gradient of chemical concentration $v$. As is well-known, the system (2) exhibits a competitive mechanism between diffusion and aggregation induced by chemotaxis where blowup—the extreme form of aggregation—can occur in finite time under certain conditions:
  • In one dimension, diffusion dominates aggregation, and all solutions of (2) are globally bounded [2].
@@
Recently, chemotaxis models involving indirect signal production, often expressed in the following typical form
$\begin{cases}u_{t}=\Delta u-\nabla\cdot(u\nabla v)+f(u,w),&x\in\Omega,t>0,\\ v_{t}=\Delta v-v+w,&x\in\Omega,t>0,\\ w_{t}=\tau\Delta w-w+u,&x\in\Omega,t>0,\end{cases}$ (3)
@@
When the signal producer is static, i.e., $\tau=0$, the indirect attractive mechanism has an explosion-suppressing (or boundedness-supporting) effect. Hu and Tao [9] showed that quadratic logistic dampening $f(u,w)=\mu u(1-u)$ to any weak extent $\mu>0$ can prevent blowup in a three-dimensional indirect chemotaxis model with growth (3). In contrast, Winkler [10] established boundedness in the three- and higher-dimensional parabolic-parabolic chemotaxis system with a strong logistic source in the sense of $\mu>\mu_{\star}$ for some $\mu_{\star}>0$. For the mass-preserving system (3) with $f\equiv 0$, a novel two-dimensional critical mass phenomenon has been identified, concerning infinite-time blowup as documented by Tao and Winkler [11] and Laurençot [12]. For chemotaxis models (3) with phenotype switching f=u+wf=-u+w, the occurrence of two-dimensional infinite-time blowup depends on the total mass $\int_{\Omega}u_{0}+\int_{\Omega}v_{0}$ [13].
@@
For fully parabolic indirect-chemotaxis models ($\tau=1$) without a source term $f(u,w)\equiv 0$, the indirect attractant mechanism can also reduce explosion-supporting potential. Fujie and Senba [14, 15] established global boundedness in physical spaces $n\leq 3$ and identified a four-dimensional critical mass phenomenon for classical solutions of (3). Specifically, they found that
@@
  • Under symmetry assumptions, the solutions exists globally and remain bounded for nonnegative initial data $(u_{0},v_{0},w_{0})$ with $\int_{\Omega}u_{0}<64\pi^{2}$. Without radial symmetry, the same conclusion holds for (3) under mixed boundary conditions, where $\partial_{\nu}u-u\partial_{\nu}v=v=w=0$ on $\partial\Omega$ for $t>0$.
@@
  • For any $m\in(64\pi^{2},\infty)\setminus 64\pi^{2}\mathbb{N}$, there exist initial data satisfying $\int_{\Omega}u_{0}=m$ such that the solutions of (3) under mixed boundary conditions blow up in finite or infinite time.
@@
Beyond four dimensions, Mao and Li [16] showed that for any $m>0$, there exist initial data with $\int_{\Omega}u_{0}=m$ such that the solutions of the system (3) blow up in finite or infinite time with symmetry assumptions. We emphasize both critical dimension $4$ and critical mass $64\pi^{2}$ of (3) regarding blowup are significantly larger compared to those of the classical chemotaxis model (2). This increase is due to the strong parabolic smoothing effect of indirect attractant production. Chen and Xiang [17] observed a similar positive effect on boundedness in certain indirect chemotaxis-haptotaxis model. Additionally, taxis-driven aggregation, particularly the spontaneous emergence of large densities during the intermediate transition triggered by large phenotype switching rates, has been explored in [18, 19].
@@
Concerning finite-time singularity formation, Tao and Winkler [20] considered a Jäger-Luckhaus variant of the system (3), where the second equation was modified to $\Delta v-\int_{\Omega}w/|\Omega|+w=0$, and the source term was taken as f=au+bwf=-au+bw for some constants $a\geq 0$ and $b\geq 0$. They showed that there exist radially symmetric initial data such that solutions of the variant, posed in a ball of five and higher dimensions, blow up in finite time. This was deduced by robust comparison methods, which are applicable to cooperative systems involving volume-filling and phenotype switching. In this regard, we focus on detecting finite-time blowup for (1), a Nagai-type variant of (3). Here, we adopt the assumption that the change of $w$ is stationary over time [21]. Notably, the system (3) is also explained as the balanced cases between attraction and repulsion in competing chemotaxis that models the aggregation of microglia observed in Alzheimer’s disease  [22, 14, 15, 23].
@@
The system (1) has an energy dissipation structure identified by Fujie and Senba [14]. The Lyapunov functional
@@
along the reasonably regular solution curves of the system (1). Here, the dissipation rate reads as
@@
$\begin{cases}-\Delta w+w=u,&x\in\Omega,\\ \partial_{\nu}w=0,&x\in\partial\Omega.\end{cases}$ (4)
Invoking the above mathematical structure, we are inspired by previous works (see Remark Remark below) to derive a lower bound for $\mathcal{F}$ in terms of a sublinear power of $\mathcal{D}$ in the form of
[…]
for a large class of radially symmetric function pairs $(u,v)=(u(r),v(r))$, which, in particular, includes the radial trajectories of the system (1). This result leads to a superlinear differential inequality for -$\mathcal{F}$, which, in turn, implies occurrence of finite-time blowup for initial data with low energy.
@@
$0\lneqq u_{0}\in C^{0}(\overline{\Omega}),\quad 0\leq v_{0}\in C^{2}(\overline{\Omega})\quad\text{with }\partial_{\nu}v_{0}\equiv 0\text{ on }\partial\Omega$ (6)
@@
[…]
@@
$\text{ such that }\|u_{0}\|_{L^{1}(\Omega)}=m,\|v_{0}\|_{W^{2,2}(\Omega)}<A\text{ and }\mathcal{F}(u_{0},v_{0})<-K(m,A)\},$
@@
Concerning the existence of initial data with low energy, for any given positive initial datum with suitable regularity, we construct a sequence of initial data with arbitrarily large negative energy, which can approximate the given initial datum in an appropriate topology.
@@
Assume $\Omega=B_{R}\subset\mathbb{R}^{n}$ for some $n\geq 5$ and $R>0$. If $u_{0}\in C^{0}(\overline{\Omega})$ and $v_{0}\in C^{2}(\overline{\Omega})$ comply with (6) and (7), then one can find a sequence
$\{(u_{k},v_{k})\in L^{\infty}(\Omega)\times W^{2,\infty}(\Omega)\}_{k\in\mathbb{N}}$
of function couples with properties: $(u_{k},v_{k})$ satisfying (6), (7) and $\int_{\Omega}u_{k}=\int_{\Omega}u_{0}$ for all $k\in\mathbb{N}$ is such that
$\|u_{k}-u_{0}\|_{L^{p}(\Omega)}+\|v_{k}-v_{0}\|_{W^{2,2}(\Omega)}\to 0\quad\text{as }k\to\infty,$
holds for all $p\in\Big[1,\frac{2n}{n+4}\Big)$, but $\mathcal{F}(u_{k},v_{k})\to$-$\infty$ as $k\to\infty$.
Remark.
Our methods are built on the original ideas of Winkler [6], and are also inspired by previous works where functional inequalities of the type in (5) have been successfully applied to chemotaxis models in various contexts: volume-filling by Cieślak and Stinner [24, 25, 26] and by Cao and Fuest [27]; the Cauchy problem by Winkler [28]; competing chemotaxis by Lankeit [29]; two species models by Li and Li [30]. As far as we are aware, this work presents the first systematic implementation of the functional inequality (5) within the context of indirect chemotaxis.
The primary challenge lies in deriving the functional inequality (5) due to explosion-suppressing effect of indirect attractant mechanisms. Mathematically, we shall obtain interior estimates $\|\Delta v\|_{L^{2}(B_{\rho})}$ in terms of the dissipation rate $\mathcal{D}$ for $\rho\in(0,R)$. To achieve this, we first establish weighted uniform estimates $\||x|^{\beta-1}v\|_{W^{1,\infty}(\Omega)}$ with some $\beta>n-2$ in Section 3. In Section 4, we are stimulated by [6, 31] to introduce the Pohožaev-type test function $\mathds{1}_{B_{\rho}}|x|^{n-2}x\cdot\nabla v$ to obtain the more complex second-order estimate. Here and below, $\mathds{1}_{E}$ stands for the indicator function of a Borel set $E\subseteq\mathbb{R}^{n}$.
This paper is organized as follows. In Section 2, we present some preliminaries involving the energy functional and basic estimates. Section 3 is dedicated to deriving uniform-in-time pointwise estimates for the chemical concentration. In Section 4, we establish the functional inequality (5). Section 5 is devoted to the proof of Theorem 1.1. In Section 6, we construct a sequence of initial data with arbitrarily large negative energy thus completing the proof of Theorem 1.2.
2 Preliminaries
We begin by revisiting the results on the local existence and uniqueness of classical solutions to the system (1).
@@ 2 Preliminaries
Let $n\in\mathbb{N}$ and $\Omega\subset\mathbb{R}^{n}$ be a bounded domain with smooth boundary. Assume that $(u_{0},v_{0})$ is as in (6). Then there exist $T_{\max}\in(0,\infty]$ and a unique triplet $(u,v,w)$ of nonnegative functions from $C^{0}\left(\bar{\Omega}\times\left[0,T_{\max}\right)\right)\cap C^{2,1}\left(\bar{\Omega}\times\left(0,T_{\max}\right)\right)$ solving (1) classically in $\Omega\times\left(0,T_{\max}\right)$. Moreover, $u>0$, $v>0$ and $w>0$ in $\Omega\times\left(0,T_{\max}\right)$, and
$\text{if }T_{\max}<\infty,\text{ then }\|u(\cdot,t)\|_{L^{\infty}(\Omega)}\to\infty\text{ as }t\nearrow T_{\max}.$ (9)
If in addition $\Omega=B_{R}$ for some $R>0$, and $(u_{0},v_{0})$ is a pair of radially symmetric functions, then $u$, $v$ and $w$ are all radially symmetric.
@@
The local existence, uniqueness and extensibility criterion of classical solutions to the system (1) can be established by adapting well-known fixed-point arguments in the contexts of chemotaxis models, as outlined in [32, Theorem 3.1], [22, Lemma 3.1] and [14, Section 4]. The nonnegativity follows from the maximum principle, while conservation of radial symmetry is a consequence of the uniqueness of solutions and the adequate form of the equations in (3). ∎
A Lyapunov functional associated with the system (1) was formulated in [14, Proposition 6.1].
@@
$\frac{\;\mathrm{d}}{\;\mathrm{d}t}\mathcal{F}(u(t),v(t))+\mathcal{D}(u(t),v(t))=0\quad\text{ for all }t\in(0,T_{\max})$ (10)
@@
$\begin{split}&\mathcal{F}(u,v)=\int_{\Omega}(u\log u-uv)+\frac{1}{2}\int_{\Omega}|(-\Delta+1)v|^{2}\quad\text{and}\\ &\mathcal{D}(u,v)=\int_{\Omega}\left(\left|\nabla v_{t}\right|^{2}+\left|v_{t}\right|^{2}\right)+\int_{\Omega}u|\nabla(\log u-v)|^{2}.\end{split}$ (11)
@@
Let $\Omega=B_{R}$ for some $R>0$ and $n\geq 4$. Then it holds that for any classical solution $(u,v,w)$ given as in Proposition 2.1,
$\int_{\Omega}u(x,t)=\int_{\Omega}u_{0}=\int_{\Omega}w(x,t)=:m\quad\text{for all }t\in(0,T_{\max})$ (12)
@@
$\int_{\Omega}v(x,t)\leq\max\left\{\int_{\Omega}v_{0},\int_{\Omega}u_{0}\right\}\quad\text{for all }t\in(0,T_{\max}).$ (13)
Furthermore, for each $p\in\Big[1,\frac{n}{n-2}\Big)$ and $q\in\Big[1,\frac{n}{n-1}\Big)$, there exist constants $C(p)>0$ and $C(q)>0$ such that for any choice of initial data $(u_{0},v_{0})$ with (6) and (7), the solution of (1) satisfies
$\|w\|_{W^{1,q}(\Omega)}\leq C(q)m\quad\text{for all }t\in(0,T_{\max}),$ (14)
@@
$\|v_{t}\|_{L^{p}(\Omega)}+\|w\|_{L^{p}(\Omega)}+\|v\|_{W^{2,p}(\Omega)}\leq C(p)(m+\|v_{0}\|_{W^{2,2}(\Omega)})\quad\text{for all }t\in(0,T_{\max}).$ (15)